METCRAX: Logbook Entries

METCRAX: Site Floor Messages, 22 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
163 Fri 03-Nov-2006flr teardownFloorOncley
161 Wed 01-Nov-2006flr TP01FloorOncley
155 Mon 30-Oct-2006flr TP01 never hooked upFloorOncley
143 Fri 27-Oct-2006Soloar panel shutdown on SHT sensorsFloorSemmer
135 Wed 25-Oct-2006SHT.0.5m adjustmentFloorSemmer
134 Wed 25-Oct-2006SHT 5m back on the airFloorSemmer
128 Wed 25-Oct-2006krypton kleanFloorOncley
126 Wed 25-Oct-2006TRH.sht.5m deadFloorSemmer
112 Sat 21-Oct-2006XML floor changedFloorSemmer
111 Sat 21-Oct-2006SHT output changeFloorSemmer
107 Sat 21-Oct-2006SHT dataFloorSemmer
106 Sat 21-Oct-2006SHT probesFloorSemmer
105 Fri 20-Oct-2006TRH fans swappedFloorOncley
101 Fri 20-Oct-2006sht RH calibration problemFloorOncley
85 Mon 16-Oct-2006Fixed RFD configurationFloorHorst
68 Fri 13-Oct-2006flr sensor serial numbersFloorMartin
53 Wed 11-Oct-2006configuration corrected for RFD tower at flr; dsm restartedFloorMartin
50 Mon 09-Oct-2006krypton changed at flrFloorMartin
26 Fri 06-Oct-2006etherant reset and flr restartedFloorMartin
21 Fri 06-Oct-2006radiometers cleanedFloorHoch
9 Thu 05-Oct-2006Stopped and started flr adamFloorHorst
8 Thu 05-Oct-2006Fuel cellFloorHorst

163: LOG, Site Floor, Fri 03-Nov-2006 21:06:56 MST, flr teardown
Serial numbers from yesterday's teardown of flr:	KZ970378	KZ030676
PSP.out	KZ940187
PYG.out	KZ040740
Tsurf	3577-2
Rnet-horiz	Q99258
Shadowband	PY7122
Gsoil	H993562	(appeared to be installed deeper -- about 10cm)
Tsoil	STP97003
TP01	200239
logger	floor	5,12177

0.5m	CSAT	671	SHT	1
1.5m	CSAT	537	SHT	2	KH20	1397
5m	CSAT	800	SHT	4
8.5m	CSAT	540
Bridge	J
161: Soil, Site Floor, Wed 01-Nov-2006 16:31:39 MST, flr TP01
At about 0900 yesterday, I removed the TP01 cabling from sw.
At about 1030 yesterday, I reinstalled it in flr's logger and we started getting
data for the first time.  The values appear to be consistent, so I think this
will give values that can be applied back to the 6 Oct rain event.

155: Soil, Site Floor, Mon 30-Oct-2006 16:20:31 MST, flr TP01 never hooked up
We have just discovered that the wiring to the Amp connector inside the logger
box for the TP01 is completely missing, thus we have never acquired data from
this sensor for METCRAX.  My guess is that the wiring might have been removed
from this box to fix another box during either TREX or CuPIDO and never 
replaced.  We plan to put another logger box at this site tomorrow and thus
get 2 days of good data before flr's teardown.

143: TRH, Site Floor, Fri 27-Oct-2006 10:20:30 MST, Soloar panel shutdown on SHT sensors
The solar panel units on the SHT sensors have shutdown. The .5m and 2m
shutdown on the night of 10/23 while the new unit at 5m shutdown last night.
These units were not designed for extensive outdoor use.

135: TRH, Site Floor, Wed 25-Oct-2006 14:22:24 MST, SHT.0.5m adjustment
I tried to align the SHT probe such that the sensor was perpendictular to the
air flow to see if that would improve the performance compared to the 50Y.
This took place during our routine visit to the floor (~11:30 to 13:00).

134: TRH, Site Floor, Wed 25-Oct-2006 14:20:17 MST, SHT 5m back on the air
A new solar panel/battery system was attached to SHT.5m. The sensor
is woking again.
128: KH2O, Site Floor, Wed 25-Oct-2006 14:04:51 MST, krypton klean
krypton kleaned ~1130.  MSB of counts went from 08 -> 17.

126: TRH, Site Floor, Wed 25-Oct-2006 09:14:54 MST, TRH.sht.5m dead
The 5m SHT sensor died last night. This the unit that does not have the solar
panel. So a full set of 3 AA batteries lastest about 1 week. A new solar
panel unit will be installed hopefully today.

112: Station_Visit, Site Floor, Sat 21-Oct-2006 15:54:50 MST, XML floor changed
The configuraton XML file was changed to handle the new SHT output message.
111: TRH, Site Floor, Sat 21-Oct-2006 15:52:48 MST, SHT output change
I changed the output message of the SHt probes so that we are collecting the
raw data for temperature and RH. This took place during our visit today, around

107: TRH, Site Floor, Sat 21-Oct-2006 09:31:48 MST, SHT data
The 5 minute averages compared to the TRH sensors are not as good as expected.
I believe this may be due to the wireless link. Each unit uses a radio talking
to a common master tied to the ADAM. The radios due not have a high level of
collision protection so some data messages may be bad. I saw this event during
lab tests. Therefore the 5 minute data may have bad data averaged in. I'll
need to look at the raw data.

106: TRH, Site Floor, Sat 21-Oct-2006 09:04:14 MST, SHT probes
On Wednesday, 18th, we installed 3 prototype Sensirion SHT75 temperature/
humidity probes. They are collocated with TRH sensors at .5m, 2m, and 5m
heights. The probes were mounted instide the TRH shield.
105: TRH, Site Floor, Fri 20-Oct-2006 18:59:41 MST, TRH fans swapped
We've seen that T.2m.flr is always higher than T.0.5m.flr, even in the day.
This usually is a symptom of a bad fan at 2m, that would cause it to read
high.  We inspected these fans today, but all were working (though 2m was
quieter and seemed to have slightly less flow).  Nevertheless, we decided
to swap the fans between these two sensors.  This was done by Steve S. between
1200-1300.  Now the profile changes as expected.  (Though is 0.5m could now bee
too high/low??)

He also took advantage of the opened-up sensors to check and possibly adjust
the SHT75 sensors he installed yesterday (or was it the other way around?)

101: TRH, Site Floor, Fri 20-Oct-2006 09:29:42 MST, sht RH calibration problem
The SHTs give bad (high) RH values when the temperature is below zero.  Steve
Semmer thinks this is an error in the unsigned math used by the PIC.  We'll
see if it is correctable, but in any case won't worry about it.

85: Radiometers, Site Floor, Mon 16-Oct-2006 11:37:28 MST, Fixed RFD configuration
Charlie corrected the aster and flr configurations to properly parse 
the RFD radiometer data.

68: Station_Visit, Site Floor, Fri 13-Oct-2006 17:27:21 MST, flr sensor serial numbers
Q7		Q99258
LICOR		PY10835
pygin		KZ30676
pygout		KZ40740
pyrin		KZ970378
pyrout		KZ940187
Tsoil		STP97003
HFT		H993562
TP01		200239

53: Radiometers, Site Floor, Wed 11-Oct-2006 12:35:02 MST, configuration corrected for RFD tower at flr; dsm restarted
The configuration had an incorrect scanfFormat for
the data logger from the RFD tower, which has caused
the Tpile to be put into the Rpile variable, and the
Tpile.8.5m to be skipped entirely. 

Fixed the configuration, downloaded to flr, and
restarted the dsm.

50: KH2O, Site Floor, Mon 09-Oct-2006 14:29:37 MST, krypton changed at flr
krypton sensor s/n 1101 was replaced with s/n 1397 
at approx. 1900 UTC 09 Oct. Unit 1101 was brought 
back to the base.

26: Adam, Site Floor, Fri 06-Oct-2006 17:06:40 MST, etherant reset and flr restarted
Flr etherant had been down since early this AM.
Cycling power fixed it.

The data system had been running and archiving
locally during this time.

21: Radiometers, Site Floor, Fri 06-Oct-2006 10:53:46 MST, radiometers cleaned
03.10. 16:18 cleaned 4 comp radiometers, CG4s , not net rad, noticed 4 comps to be somewhat loose. 
9: Adam, Site Floor, Thu 05-Oct-2006 10:07:12 MST, Stopped and started flr adam
I accidently started additional dsm processes on the flr adam.
Stopped and restarted adam to kill them.
8: Power, Site Floor, Thu 05-Oct-2006 09:17:51 MST, Fuel cell

Tried to install the Voler fuel cell but could not find the required stainless
steel rod required to connect the hydrogen bottle.  John and Kurt had previouslystarted to install the PowerPEM D35, but Steve prefers to use the Voler.  He 
would like it to be used to just charge spare batteries, independent of 
the solar panels.